Which of these pipes will
empty a pool the fastest?
1.One pipe of diameter 60 cm.
2.Two pipes of diameter 30 cm each.
3.Three pipes of diameter 20 cm each.
Please explain

Answer by richwmiller(17219)

Although it looks like they are equal the areas and volumes would be different.
a=pi*30^2,
b=2*pi*15^2,
c=3*pi*10^2
{a = 900 pi, b = 450 pi, c = 300 pi}
a=2827.4, b=1413.7, c=942.48
The single pipe would carry more liquid faster